The total of two numbers is 17. The larger number subtracted from three times the smaller number is 13 less than the sum of the larger number and twice the small number. Find the numbers.
Answer by checkley77(12844)

x+y=17 or x=17-y
3y-x=x+2y-13 replace x by (17-y) in this equation & solve for y.
3y-(17-y)=(17-y)+2y-13
3y-17+y=17-y+2y-13
3y+y+y-2y=17-13+17
3y=21 ans.
y+21/3
y=7 ans.
3*7-x=x+2*7-13
21-x-x=14-13
-2x=1-21
-2x=-20
x=-20/-2
x=10 ans.
Proof:
3*7-10=10+2*7-13
21-10=10+14-13
11=11